OpenStudy (anonymous):

the sum of two consecutive integers is 73

OpenStudy (solomonzelman):

Name the integers x and x+1 \(\normalsize\color{blue}{ x+(x+1)=73 }\)

OpenStudy (anonymous):

@SolomonZelman is a little bich

OpenStudy (solomonzelman):

Need more help @MelissaRuth ? (Ignore Mrking123)

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

helppppp

OpenStudy (camerondoherty):

an easy way to do it also is subtract one from it and divide by 2

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ x+(x+1)=73 }\) can you solve for x? After you solve for x, find the first number ` x ` and the second number `x+1

OpenStudy (camerondoherty):

soooo 73 - 1 = 72 72 div 2 = 36

OpenStudy (camerondoherty):

Its called working backwards